July 2, 2013 Calorimetry

From GlueXWiki
Jump to: navigation, search

'Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Light guides
    • Shipment of replacement light guides (USM)

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Monitoring
  4. Proposed dimensions for Fcal monitoring pane
  5. FCAL
  6. BCAL
  7. Reconstruction and Simulation
  8. TDR update
  9. Calibration of Calorimeters
  10. Any other business


Attendees: Elton, George, John (JLab); Christina, Pavlos (Athens); Shawn, Andrei (UofR); Will, Alam, Elias (USM).

  1. Announcements
  2. Action Items
  3. Monitoring
      • Christina: Payment not yet received. Elton will check up on it. [Note added: Christina sent a note on Wed saying the payment had been received].
      • Elton: 43 Bcal LED boards have been repaired by soldering new LEDs. 16 more boards have the soldering pad stripped and could be repaired but with more difficulty. It appears we will have sufficient boards with the ones that George brought with him.
    1. Proposed dimensions for Fcal monitoring pane
      • Chuck has made the drawings for the plexiglass panes and they have been checked by John Frye.
      • Elton has requested information from Athens regarding the polishing of the edges away from the beam. We have decided to order the plexiglass material cut to the correct dimensions but paint, sand, and polish at JLab. That will allow for some iteration with the process once we have the plexiglass.
    2. NIM paper draft has been sent out for feedback
  4. FCAL (John Leckey)
    • Last labyrinth components are planned to ship out from IU tomorrow and should arrive early next week
    • Should start cabling middle of next week. Shawn and Zach (from UofR) will help with the cabling
    • Orders for the environmental monitoring have bee placed. A prototype of the custom box with temp, humidity, and light sensors have been built and work well. They provide 0-5 V analog output signals that can be registered by the PLC. Have also ordered test crash buttons. Sufficient components for boxes plus spares have been ordered. Cables are color coded to facilitate repairs by non-experts.
    • The testing of individual signals has been completed. They continue to add complexity as they integrate and assemble the full system.
    • Have checked the power to the magnets, which will be used to lock the doors.
    • Picked out lighting for the dark room. Red lights can be on for work with bases powered. White lights are interlocked to the custom box signals though the PLC interface.
  5. BCAL
    • GlueX-doc-2248 Light transmission measurements in Bcal (George)
    • Stability of transmission measurements over a week in the ESB verified with module 10
    • Change of normalization of transmission signal over a two month period seen with module 47.
    • Christina suggested that temperature may be a variable that needs to be watched. See NIM draft for temperature response.
  6. USM
    • Light guides (Elias): Plan to ship 65 replacement guides by Friday. Will continue production of 3 full sets once these have been completed.
    • Alam is still waiting on his invitation letter.
  7. Reconstruction and Simulation
  8. TDR update
  9. Calibration of Calorimeters
  10. Any other business